Write a quadratic equation and solve
Mr. Tobias runs a hot dog distribution warehouse that distributes hot dog to local grocery stores. The amount of profit he makes depends on the number of boxes of hot dogs he sells. The profit follows the equation P(x) = -x^2 + 140x - 3940. What is his maximum profit, and how many boxes must he sell to make the maximum profit?

Answer by stanbon(75887)

Mr. Tobias runs a hot dog distribution warehouse that distributes hot dog to local grocery stores. The amount of profit he makes depends on the number of boxes of hot dogs he sells. The profit follows the equation P(x) = -x^2 + 140x - 3940. What is his maximum profit, and how many boxes must he sell to make the maximum profit?
--------------
Max occurs when x = -b/(2a) = -140/(-2) = 70 boxes
------
Cheers,
Stan H.
------------